OpenStudy (anonymous):

Find the slope of the line that passes through the points (8, 7) and (11, 5). Enter fractions as reduced improper fractions if needed, like this. 4/3

hartnn (hartnn):

apply same formula! should i mention it again ?

OpenStudy (anonymous):

oh how would i reduce it though?

OpenStudy (primeralph):

@Kittyxb3ll Is your arithmetic good?

OpenStudy (anonymous):

arithmitic?

OpenStudy (primeralph):

Yes, arithmetic.

OpenStudy (anonymous):

whats that

OpenStudy (primeralph):

Never mind.

OpenStudy (anonymous):

oh..

hartnn (hartnn):

what did u get after u applied the formula ?

OpenStudy (anonymous):

hold on let me do it real quick :0

OpenStudy (anonymous):

2 over 3

hartnn (hartnn):

+2/3 or -2/3 ??

OpenStudy (anonymous):

+ or should i have gotten a 2/-3?

hartnn (hartnn):

yes, try again ?

OpenStudy (anonymous):

oh uhm...

hartnn (hartnn):

7-5 = 2 8-11 = -3 so, m = -2/3 and it cannot be further reduced as 2 doesn't go into 3 :)

OpenStudy (anonymous):

oh okay but the -... is it for both or just the 3?

hartnn (hartnn):

what u think ?

OpenStudy (anonymous):

both? like make the whole fraction -..? right?

hartnn (hartnn):

when either numerator OR denominator is negative, the whole fraction is negative, when both are negative, the fraction becomes positive

OpenStudy (anonymous):

ohh okay I'll put that in my notes :3 thanks again

hartnn (hartnn):

welcome ^_^